• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ibrahim

    Log entry time 00:09:38 on November 7,2008

    Entry number 247190

    keyword=Swing Shift Summary

    Swing Shift Summary

    16:00 Shift started while the hall in restricted access.

    17:00 The Left Dipole polarity was changed to Negative by hall A tech.

    17:40 The Left arm quads polarity was changed to Negative by the RC.

    17:45 MCC are doing the hall sweep.

    17:50 Left arm momentum is set to 2.4 GeV/c.

    18:40 Hall in controlled access.

    19:35 Moved the target to the Empty position

    19:40 Hall to power permit.

    20:40 Tune beam to the hall.

    20:58 MCC still need few more minutes to check something wrong with the beam.

    10:00 The Hall_A HV alarm goes off every 1 or minutes and clears itself right away. I asked Alexander about this and he said it is probably not serious fluctuations.

    11:00 A new set of beam positions are in place:

    For He3 Cell: BPMA = (1,1) and BPMB = (-1,1)

    For all other targets: BPMA = (1,4) and BPMB = (-1,4)

    11:10 The MCC are still having problems with the beam.

    00:00 Shift ends while we are still for the beam.


    ******************************************